The fromula A=P(1+r)^t gives the amount A in dollars that P dollars will grow to in two years at interest rate r using compound interest. What interest rate will cause $6235.75 to grow to $8612.88 in two years? Round your answers here to the nearest tenth of a percent.: The fromula A=P(1+r)^t gives the amount A in dollars that P dollars will grow to in two years at interest rate r using compound interest. What interest rate will cause $6235.75 to grow to $8612.88 in two years? Round your answers here to the nearest tenth of a percent.
Answer by nerdybill(465)

The fromula A=P(1+r)^t gives the amount A in dollars that P dollars will grow to in two years at interest rate r using compound interest. What interest rate will cause $6235.75 to grow to $8612.88 in two years? Round your answers here to the nearest tenth of a percent.
.
The problem gives a formula of:
A=P(1+r)^t
.
It also provides:
P = $6235.75
A = $8612.88
t = 2 years
.
Simply plug in the above into the provided formula and solve for unknown (r):
A=P(1+r)^t
8612.88 = 6235.75(1+r)^2
dividing both sides by 6235.75 we get:
1.3812 = (1+r)^2
take the "square root" of both sides to get:
1.1752 = 1+r
subtract 1 from both sides to get:
.1752 = r
.
answer: 17.5%